OpenStudy (anonymous):

I know this is not the correct term but how do I reverse a square root. For example 15 squared is 225. If I given 225 how can i find what I used to find 225

OpenStudy (anonymous):

You take square root. \[\sqrt{225}=15\]

OpenStudy (anonymous):

In general, \[\sqrt{a^2}=(\sqrt{a})^2=a\]